A random sample of size 20 drawn from a normal distribution has a sample mean of 8 and a sample standard deviation of 4.
A) Compute the sample test statistic t under the null hypothesis H_0: Mue(mean)=7
B) State the alternate hypothesis if you believe that the population mean may be more than 7.

Answer by stanbon(75887)

A random sample of size 20 drawn from a normal distribution has a sample mean of 8 and a sample standard deviation of 4.
A) Compute the sample test statistic t under the null hypothesis
Ho: Mue(mean) = 7
t(8) = (8-7)/[4/sqrt(20)] = 1/0.8944 = 1.118
---------------------------------------------
B) State the alternate hypothesis if you believe that the population mean may be more than 7.
Ha: u > 7
---------------------
Cheers,
Stan H.
-------------